# DAWN: Domain-Adaptive Weakly Supervised Nuclei Segmentation via Cross-Task Interactions

### v1

@@ IV-E4 Qualitative Comparisons
We provided visualization comparisons of different methods in Fig. 7 and Fig. 8. From Fig. 7, comparing with only using source domain method, our domain adaptation method exhibits improved recognition capability for overlapping and missed nuclei by incorporating an output of detection network in CPL. In Fig. 8, our results are closest to the nuclei segmentation ground truth. In contrast, SIFA[17] and PDAM[18] exhibit poorer segmentation performance on some overlapping nuclei. These results further demonstrate the effectiveness of our proposed method in domain adaptation, whether from pan-cancer to specific cancer types or from pan-cancer to other pan-cancer datasets.
TABLE VI: The ablation experiments of CFC and IST modules on TNBC, CryoNuSeg, Lizard and ConSeP datasets.
@@ IV-F Ablation study
IV-F1 Evaluation of CFC and IST
To prove the effectiveness of CFC and IST components, we perform ablation studies with results listed in Table VI. On the TNBC and Lizard datasets, the five metrics have noticeable improvements whether CFC or IST is added. On CryoNuSeg and ConSeP datasets, when the CFC module is used, the DQ and PQ also obtain improvement, and the DQ and AJI decrease slightly. At the same time, when adding CFC and IST components, the model achieves the best performance, indicating that CFC and IST modules need to be used together. This phenomenon can be attributed to the significant morphological differences between the CryoNuSeg (ConSeP) and the source domain datasets. Only using CFC might lead to feature perturbations of the segmentation network. However, adding IST can guide the network in performing domain transfer toward the target domain. Our ablation study proves that the CFC and IST promote domain adaptation ability and enhance nuclei segmentation performance on weakly supervised scenarios.
TABLE VII: The mask fusion and distance filter ablation experiments on TNBC dataset.
@@ IV-F1 Evaluation of CFC and IST
ours(d=10) 0.6541 0.4917 0.5771 0.7054 0.4170
ours(d=25) 0.8561 0.7280 0.8425 0.8202 0.6925
ours(d=40) 0.8332 0.7120 0.8313 0.8220 0.6850
@@
TABLE VIII: The detection results before and after domain adaptation.
Datasets Stage Recall Precision F1
TNBC Bef. 0.7647 0.8949 0.8205
CryoNuSeg Bef. 0.4637 0.7522 0.5737
Lizard Bef. 0.7804 0.8373 0.7958
ConSeP Bef. 0.5563 0.7893 0.6526
IV-F2 Evaluation of Hyper-parameter in CPL
To prove that mask fusion and distance filter effectively improve segmentation performance, we perform an ablation study and parameter-sensitive experiments on the TNBC dataset. The results are shown in Table VII. From the result, we can see that, whether removing mask fusion or distance filter, the performance decreases evidently, which indicates the two components are effective in the CPL module because they can enhance the quality of pseudo-labels. Besides, we can see that the model is sensitive to the distance filter method. When the distance is small, the performance decreases because it destroys the nuclear shape. When the distance is enlarged, the performance is close to the result without a distance filter. Overall, the mask fusion and distance filter in CPL have also been proved to improve performance.
IV-F3 Evaluation of Iterative Training
We perform the experiments across the four datasets to evaluate the impact of the number of iterations in Fig. 9. In this figure, the x-axis represents the number of iterations, corresponding to the parameter ”r” in Algorithm 1, while the y-axis represents the values of the DICE, AJI, and PQ metrics. The results show that on the TNBC and Lizard datasets, the model’s performance shows an increasing trend as the number of training iterations increases. Specifically, the TNBC dataset achieves optimal performance at r=3, while the Lizard dataset performs best at r=4. In contrast, on the CryoNuSeg and ConSeP datasets, the model’s performance remains relatively stable after the second round of training. Consequently, we select the optimal performance according to r as the segmentation result. Furthermore, Fig. 10 presents the loss function and F1 curve for three iterations of training on the TNBC dataset, with each iteration consisting of 20 training epochs. From this figure, it can be observed that as training progresses, the model’s loss steadily decreases. Correspondingly, the F1 score increases with training and performs best in the final iteration.
IV-F4 Evaluation of Domain Adaptation
To evaluate the influence of our proposed auxiliary detection branch on nuclei detection ability, we conduct a domain adaptation comparison experiment on four target datasets in three metrics, i.e., Recall, Precision, and F1. The experimental results are shown in Table VIII. As seen from the table, when using the detection branch to assist domain transfer, the Recall has been significantly improved without reducing F1, which means that the auxiliary task can help the segmentation network to identify the nucleus more completely. This conclusion can also be verified from the visualization results in Fig. 11. The figure’s red $\star$ represents the actual nuclei location. Orange $\bullet$” and blue “+” represent the predicted location before and after domain adaptation. The figure shows that after adding the detection task, the predicted results are highly coincident with the actual labels, indicating that our detection task helps enhance the nuclei recognition ability of the segmentation task.
V Conclusion
@@
To reduce the dependence of nuclei segmentation on manual annotation, we propose a domain-adaptive weakly supervised segmentation method named DAWN. Our method can transfer the model trained on the public data (source domain) to arbitrary external data (target domain) without fine pixel-level annotation. For this purpose, the domain adaptation of the segmentation network is guided by an auxiliary detection task. In DAWN, a consistent feature constraint (CFC) module is used to enhance the domain transfer efficiency at the feature level. In addition, a combined pseudo-labels optimization (CPL) method and interactive supervision training (IST) method are proposed to enhance the domain adaptation ability at the task level. The ablation experiments prove these components are effective in boosting the nuclei segmentation performance. Furthermore, our comparison experiments show that our DAWN outperforms existing weakly supervised methods and achieves similar performance compared with supervised methods.

### v3

@@ IV-E4 Qualitative Comparisons
We provide visualization comparisons of segmentation results between domain adaptation methods in Fig. 7. First, our model exhibits better nuclei segmentation capability compared to the source domain method, especially for missed and overlapping nuclei. In contrast, SIFA[19] and PDAM[20] exhibit poor segmentation performance on some overlapping nuclei. These results further demonstrate the effectiveness of our proposed method in domain adaptation, whether from pan-cancer to specific cancer types or from pan-cancer to other pan-cancer datasets.
IV-E5 Detection performance
To evaluate the influence of our proposed auxiliary detection branch on nuclei detection, we conduct a domain adaptation comparison experiment on four target datasets in three metrics, i.e., Recall, Precision, and F1. The experimental results are shown in Table VI. As seen from the table, when using the detection branch to assist domain transfer, the Recall has been significantly improved without reducing F1, which means that the auxiliary task can help the segmentation network to identify the nucleus more completely. This conclusion can also be verified from the visualization results in Fig. 8. The figure’s red $\star$ represents the actual nuclei location. Orange $\bullet$” and blue “+” represent the predicted location before and after domain adaptation. The figure shows that after adding the detection task, the predicted results are highly coincident with the actual labels, indicating that our detection task helps enhance the nuclei recognition ability of the segmentation task.
TABLE VI: The detection results before and after domain adaptation.
Datasets Stage Recall Precision F1
TNBC Bef. 0.7647 0.8949 0.8205
CryoNuSeg Bef. 0.4637 0.7522 0.5737
Lizard Bef. 0.7804 0.8373 0.7958
ConSeP Bef. 0.5563 0.7893 0.6526
TABLE VII: The ablation experiments of CFC and IST modules on TNBC, CryoNuSeg, Lizard and ConSeP datasets.
@@ IV-F Ablation study
IV-F1 Evaluation of module designs
To prove the effectiveness of CFC and IST components, we perform ablation studies with results listed in Table VII. On the TNBC and Lizard datasets, the five metrics have noticeable improvements whether CFC or IST is added. On CryoNuSeg and ConSeP datasets, when the CFC module is used, the DQ and PQ also obtain improvement, and the DQ and AJI decrease slightly. At the same time, when adding CFC and IST components, the model achieves the best performance, indicating that CFC and IST modules need to be used together. This phenomenon can be attributed to the significant morphological differences between the CryoNuSeg (ConSeP) and the source domain datasets. Only using CFC might lead to feature perturbations of the segmentation network. However, adding IST can guide the network in performing domain transfer toward the target domain. Our ablation study proves that the CFC and IST promote domain adaptation ability and enhance nuclei segmentation performance on weakly supervised scenarios.
To prove that mask fusion and distance filter effectively improve segmentation performance, we perform an ablation study and parameter-sensitive experiments on the TNBC dataset. The results are shown in Table VIII. From the result, we can see that, whether removing mask fusion or distance filter, the performance decreases evidently, which indicates the two components are effective in the CPL module because they can enhance the quality of pseudo-labels.
TABLE VIII: The mask fusion and distance filter ablation experiments on TNBC dataset.
@@ IV-F1 Evaluation of module designs
ours 0.8561 0.7280 0.8425 0.8202 0.6925
TABLE IX: The hyper-parameter $r_{1}$ ablation experiments on TNBC dataset.
Methods PanNuke $\rightarrow$ TNBC
DICE AJI DQ SQ PQ
$r_{1}$=5 0.8147 0.6828 0.7914 0.7784 0.6554
$r_{1}$=11 0.8561 0.7280 0.8425 0.8202 0.6925
$r_{1}$=20 0.8323 0.7024 0.8191 0.7923 0.7022
IV-F2 Evaluation of Hyper-parameter
First, by observing Table IX, there is a significant decrease in the performance of the model when the radius range of the nuclei is altered. Whether $r_{1}$ is set to $5$ or $20$, the Dice coefficient drops by more than 2 %. This indicates that the setting of nucleus radius has a substantial impact on the model and that configuring parameters based on the radius from the validation set is advantageous. Secondly, observing Tables X and XI it can be seen that there is no significant variation in model performance when modifying the standard deviation $\sigma$ and $\theta$ of the nuclear radius. In fact, in some metrics, the model even outperforms the reported results in the main text; for example, when $\sigma$ = 4, the SQ metric improves by 0.2%. However, this variation is not substantial, indicating that the model parameters are less affected by $\sigma$ and $\theta$. Lastly, we also include additional sensitivity analyses of distance $d$ as presented in Table XII.
IV-F3 Evaluation of Iterative Training
We perform the experiments across the four datasets to evaluate the impact of the number of iterations in Fig. 9. In this figure, the x-axis represents the number of iterations, corresponding to the parameter ”I” in Algorithm 1, while the y-axis represents the values of the DICE, AJI, and PQ metrics. The results show that on the TNBC and Lizard datasets, the model’s performance shows an increasing trend as the number of training iterations increases. Specifically, the TNBC dataset achieves optimal performance at I=3, while the Lizard dataset performs best at I=4. In contrast, on the CryoNuSeg and ConSeP datasets, the model’s performance remains relatively stable after the second round of training. Consequently, we select the optimal performance according to I as the segmentation result. Furthermore, Fig. 10 presents the loss function and the F1 curve for three training iterations in the TNBC dataset, each iteration consisting of 20 training epochs. From this figure, it can be observed that as training progresses, the model’s loss steadily decreases. Correspondingly, the F1 score increases with training and performs best in the final iteration.
TABLE X: The hyper-parameter $\sigma$ ablation experiments on TNBC dataset.
Methods PanNuke $\rightarrow$ TNBC
DICE AJI DQ SQ PQ
$\sigma$=2.0 0.8447 0.7120 0.8394 0.8191 0.7022
$\sigma$=2.75 0.8561 0.7280 0.8425 0.8202 0.6925
$\sigma$=4.0 0.8402 0.7221 0.8334 0.8227 0.6896
TABLE XI: The hyper-parameter $\theta$ ablation experiments on TNBC dataset.
Methods PanNuke $\rightarrow$ TNBC
DICE AJI DQ SQ PQ
$\theta$=0.1 0.8489 0.7233 0.8510 0.8212 0.6917
$\theta$=0.2 0.8561 0.7280 0.8425 0.8202 0.6925
$\theta$=0.3 0.8550 0.7226 0.8390 0.8226 0.6898
TABLE XII: The hyper-parameter $d$ ablation experiments on TNBC dataset.
Methods PanNuke $\rightarrow$ TNBC
DICE AJI DQ SQ PQ
$d$=10 0.6541 0.4917 0.5771 0.7054 0.4170
$d$=25 0.8561 0.7280 0.8425 0.8202 0.6925
$d$=40 0.8332 0.7120 0.8313 0.8220 0.6850
@@ IV-F3 Evaluation of Iterative Training
Figure 11: The limitation of our proposed DAWN when facing nuclei with significant morphological variations.
V Conclusion and Limitation
@@
In this paper, to reduce the reliance on manual annotations for nuclei segmentation, we propose a domain-adaptive weakly supervised segmentation method, termed DAWN. Our approach enables the transfer of a model trained on public data (source domain) to arbitrary external datasets (target domain) without requiring fine pixel-level annotations. This domain adaptation is guided by an auxiliary detection task. DAWN employs a Consistent Feature Constraint (CFC) module to improve domain transfer efficiency at the feature level. Additionally, we introduce a combined pseudo-label optimization method (CPL) and an interactive supervision training (IST) strategy to enhance domain adaptation at the task level. Ablation experiments confirm the effectiveness of these components in improving nuclei segmentation performance.
Despite these advancements, DAWN exhibits limitations when handling nuclei with significant intra-image morphological variations, such as large spindle-shaped nuclei coexisting with small, circular ones, as illustrated in Fig. 11. This leads to occasional segmentation errors, particularly under-segmentation of nuclei with atypical shapes. To address these challenges, future work will focus on developing enhanced interaction mechanisms between the segmentation and detection networks. Specifically, we aim to introduce a multi-task learning framework that simultaneously optimizes segmentation and shape detection, allowing the model to better differentiate and capture varying nuclear shapes by leveraging contextual and spatial information across scales. Additionally, we plan to incorporate a region-aware feature aggregation module, which will allow the model to adaptively focus on regions with high morphological diversity. This approach aims to improve segmentation accuracy for nuclei with atypical geometries by explicitly modeling the relationships between various shapes within the same image. By combining these stratesgies, we aim to improve DAWN’s ability to segment diverse and complex nuclei, pushing the boundaries of domain adaptation in histopathology image analysis.